# CaSTFormer: Causal Spatio‐Temporal Transformer for Driving Intention Prediction

### v1

@@ Abstract
Accurate prediction of driving intention is key to enhancing the safety and interactive efficiency of human-machine co-driving systems. It serves as a cornerstone for achieving high-level autonomous driving. However, current approaches remain inadequate for accurately modeling the complex spatio-temporal interdependencies and the unpredictable variability of human driving behavior. To address these challenges, we propose CaSTFormer, a Causal Spatio-Temporal Transformer to explicitly model causal interactions between driver behavior and environmental context for robust intention prediction. Specifically, CaSTFormer introduces a novel Reciprocal Shift Fusion (RSF) mechanism for precise temporal alignment of internal and external feature streams, a Causal Pattern Extraction (CPE) module that systematically eliminates spurious correlations to reveal authentic causal dependencies, and an innovative Feature Synthesis Network (FSN) that adaptively synthesizes these purified representations into coherent spatio-temporal inferences. We evaluate the proposed CaSTFormer on the public Brain4Cars dataset, and it achieves state-of-the-art performance. It effectively captures complex causal spatio-temporal dependencies and enhances both the accuracy and transparency of driving intention prediction.
@@ Introduction
Driver intention prediction is critical in autonomous driving systems, playing a significant role in effectively mitigating risks and enhancing driving safety. By forecasting potential outcomes several seconds in advance, the system can alert the driver proactively or initiate evasive maneuvers, further increasing its safety capabilities.
@@
Typically, the intention prediction relied on the extraction and fusion of visual features to perform simple intention prediction (Huang et al. 2022). The development of sensor technology has enabled the use of multi-modal information, including GPS coordinates, vehicle speed, map data, and driver head pose (Hu et al. 2021; Mo et al. 2023; Li, Zhao, and Wang 2022; Wu et al. 2023). By leveraging complex models for feature extraction, fusion, and prediction, the performance of driver intention prediction has been significantly improved (Sui et al. 2021; Guo et al. 2023; Liu, Wu, and Wang 2023; Gao et al. 2023).
@@
Figure 1: Comparison between previous driving intention prediction methods and ours. (a) is an LSTM–GRU–based framework in which internal and external streams are processed independently and then concatenated for final fusion. (b) is our CaSTFormer, a Transformer–based model that performs dynamic causal fusion of the dual streams while integrating intention priors. By jointly capturing global and local dependencies and leveraging intrinsic cross-stream interactions, our method surpasses existing approaches.
@@
Despite the progress made in driver intention prediction, the rich multi-modal data has not been fully exploited. Most existing work merely concatenates or linearly aggregates this information, as illustrated in Fig. 1 (a). However, as the controller of the vehicle, changes in the driver’s state subsequently influence the vehicle’s driving status, revealing a tight dependency. Therefore, we propose to explicitly model the causal relationship between the driver and the environment, emphasizing the decisive impact of this causality on the prediction task, as shown in Fig 1 (b).
@@
Specifically, we adopt a transformer-based architecture and introduce three sequential components to enhance the causal modeling, encode the driver’s intention, and fuse the multi-dimensional features. First, we introduce a Reciprocal Shift Fusion (RSF) module that cross-fuses internal and external features through a shifting mechanism, explicitly establishing a temporal dependency between the two feature streams. As the fused features may contain considerable casual noise, we further devise a Causal Pattern Extraction (CPE) module to filter out such noise to obtain a clearer causal representation. Finally, we utilize a Feature Synthesis Network (FSN) that employs a gating mechanism to integrate the in-cabin, out-of-cabin, and interaction representations, enabling the modeling of both local and global causal structures. Our main contributions are as follows:
@@
  • We propose CaSTFormer, an efficient Transformer-based framework for driving-intent prediction that integrates temporal dependency modeling, causal pattern enrichment, and adaptive multi-view fusion within a unified end-to-end architecture.
  • CaSTFormer disentangles spatio‐temporal interactions of interior and exterior streams, subsequently isolates genuine causal contributions through counterfactual attention subtraction and adaptively fuses heterogeneous features, thereby mitigating spurious correlations and improving robustness under complex driving conditions.
@@
  • Extensive evaluation on the Brain4Cars dataset demonstrates that CaSTFormer demonstrates superior performance in driving intention prediction in both highway and urban scenarios.
@@
In the early stages of research on driving intention prediction, some studies focused on learning spatio-temporal features directly from raw video data. For instance, a combination of 3D Convolutional Neural Networks and Long Short-Term Memory networks was utilized to predict driving maneuvers directly from video streams (Gebert et al. 2019). To achieve a more comprehensive basis for decision-making, multi-modal data was introduced to improve the accuracy of intention prediction. In one such study, the in-cabin driver’s view was combined with the external traffic scene, using a Convolutional Long Short-Term Memory network to fuse and analyze features from both views (Rong, Akata, and Kasneci 2020). However, their contextual capacity is limited, making it difficult to process long-range, non-consecutive events.
@@
Inspired by human cognitive processes, TIFN (Guo et al. 2023) proposed a State Update Unit to incorporate the influence of environmental information into driver state modeling, and extracted semantic segmentation features to provide clear cues affecting driver attention. Meanwhile, another approach treated intention prediction as a sequence labeling task, further capturing the contextual dependencies of driving behavior by combining a Bidirectional Long Short-Term Memory network with a Conditional Random Field (Zhou et al. 2021). Although these models fuse multi-source information, they typically learn the correlations between features implicitly, without explicitly decoupling and reasoning about the mutual influences between different information modalities.
Furthermore, personalized differences among drivers were also taken into consideration as a significant source of model error. Early work used a Domain-Adversarial Recurrent Neural Network (DA-RNN) to learn driver-agnostic, domain-invariant features to enhance the model’s generalization capabilities (Tonutti et al. 2019). Recently, IDIPN (Liu et al. 2025) utilized Inverse Reinforcement Learning to extract unique driving preferences from each driver’s historical data, thereby correcting the predictions of a general-purpose model, while FedPRM (Zhu et al. 2024) achieved this by using a federated learning framework to directly construct personalized models for each driver.
@@
The Transformer architecture, with its unique attention mechanism, excels at capturing long-range dependencies and global interaction information in temporal data (Liu et al. 2024). CemFormer (Ma et al. 2023) integrates data from both in-cabin and external cameras, learning a unified cross-view representation via a spatio-temporal tansformer to infer driver intention directly from their behavior. However, such methods primarily rely on the inherent structure of the transformer to interpret the dependencies. Meanwhile, a non-autoregressive transformer with hybrid attention has been utilized to simultaneously capture the temporal dynamics of a single vehicle, as well as the interactions among multiple vehicles (Jiang et al. 2024). DriveTransformer (Jia et al. 2025) further constructs a unified end-to-end framework to handle perception, prediction, and planning tasks in parallel. This treats intention prediction as an integrated component of the overall scalable system, but may dilute the model’s focus on driver-intention features.
@@
Besides, some early works have attempted to introduce causal inference methods into the intention prediction task. One study constructed a causal model to learn the temporal relationships of invariant representations from driving data, thereby achieving domain generalization (Hu et al. 2022). Another adopted a driver-centric approach, formulating the task of risk object identification as a causal problem and proposing a two-stage framework based on causal inference (Li, Chan, and Chen 2020). However, the scope of scenarios and conditions covered by these methods is relatively limited.
@@
Apart from all the prior work, our method employs a dual-stream transformer. Building upon an encoding of driver intention, it explicitly establishes both local and global causal dependencies between in-cabin and external factors. This approach, through the fusion of multi-dimensional features, enables precise intention prediction to address increasingly complex driving scenarios.
@@
Figure 2: Overview of the CaSTFormer pipeline. After data preprocessing, external optical flow is encoded by ResNet-18 and interior images by MobileFaceNet to produce dual-stream feature sequences. These are then fed into three core modules: (a) Reciprocal Shift Fusion (RSF) for temporal feature integration; (b) Causal Pattern Extraction (CPE) for causal enhancement and intent embedding; and (c) Feature Synthesis Network (FSN) for dynamic fusion of complementary internal, external and interaction views to yield the final driving intention prediction.
@@
As illustrated in Fig. 2, CaSTFormer takes as input a bi-stream image sequence =$\mathcal{I}=\{(I^{\mathrm{out}}_{b,t},I^{\mathrm{in}}_{b,t})\}_{b,t=1}^{B,T}$ comprising $B$ samples, each consisting of $T$ frames from two synchronized streams (outside and inside views), where $b$ and $t$ index the sample and temporal frame, respectively. Feature extractors separately encode each stream into features $F^{\mathrm{out}},F^{\mathrm{in}}\in\mathbb{R}^{B\times T\times D}$ ($D$ denotes the dimensionality of each encoded feature vector). The obtained features pass through a Reciprocal Shift Fusion (RSF) module for temporal causality modeling via our proposed bi-stream attention, followed by a Causal Pattern Extraction (CPE) module to inject learnable causal representations. Finally, visual features ($z_{\mathrm{in}},z_{\mathrm{out}}$ and interaction features $z_{\mathrm{ctx}}$) are adaptively fused by a Feature Synthesis Network (FSN) into a joint prediction $\ell_{\mathrm{joint}}$.
Reciprocal Shift Fusion (RSF)
On multi‐lane highways, cameras concurrently record the external traffic scene and the driver’s internal state. Recognizing the mutual influences between environmental context and driver behavior, our dual‐stream architecture explicitly captures their bidirectional interactions by jointly modeling both feature streams. To enforce exterior-first precedence, we introduce a temporal delay mechanism in the Key and Value sequences. Specifically, at time step $t$, the attention mechanism accesses only information from the preceding frame $t-1$. Concretely, we define the delayed feature
[…]
@@
Under a strict single-frame delay constraint, BDA enriches each frame’s representation by fusing internal and external contexts from the immediately preceding timestep. The current internal and external features attend bidirectionally to their one-frame–delayed counterparts, capturing both temporal coherence and cross-stream coupling. To model diverse associations efficiently, we project into $H$ parallel attention heads and aggregate their outputs through concatenation and a final linear mapping:
@@
where $[\cdot]_{i=1}^{H}$ denotes head-wise concatenation and $W^{O}$ restores the original feature dimension. Figure 3 illustrates the detailed bidirectional query–key–value fusion, showing how the internal and external streams are jointly refined.
Figure 3: Illustration of Bidirectional Dependency Attention (BDA) where buffered internal and external features are mutually cross-attended to produce dynamically enhanced current-frame representations.
@@
Although BDA produces a richly fused representation $H_{b,t}$ at each spatial–temporal location, individual channels may still carry irrelevant or noisy information. Hence, we employ two successive channel‐wise gating layers that adaptively emphasize informative channels and suppress spurious responses:
$g_{b,t}$ $=\sigma\bigl{(}W_{2}\,(\mathrm{ReLU}(W_{1}H_{b,t}+b_{1}))+b_{2}\bigr{)}$ (3)
$\widetilde{H}_{b,t}$ $=g_{b,t}\,\odot\,H_{b,t}$
where $\odot$ denotes the Hadamard product between vectors.
@@
$R(x)=x\;\odot\;\frac{s}{\sqrt{\frac{1}{D}\sum_{d=1}^{D}x_{d}^{2}+\epsilon}}$ (4)
${X}_{b,t}=\mathrm{Dropout}\bigl{(}R(\widetilde{H}_{b,t})\bigr{)}$ (5)
where $R(\cdot)$ denotes the Root-mean-square normalization function, $s$ is a learnable scaling parameter, $\epsilon$ is a small constant to ensure numerical stability, and $D$ is the dimensionality of the feature vector $x$, yielding channel-calibrated representations across both internal and external feature streams.
Causal Pattern Extraction (CPE)
Conventional intention prediction architectures aggregate heterogeneous interior and exterior cues under an implicit correlation assumption and thus often mistake coincidental patterns for true decision drivers. To overcome this limitation, our CPE module contrasts observed and counterfactual cross‐stream attentions to disentangle direct causal contributions, then selectively amplifies only those residuals that genuinely influence driving intent, resulting in improved robustness and generalization in safety‐critical scenarios. Specifically, CPE takes the bidirectionally fused interior and exterior features ${X}_{b,T}^{\mathrm{in}},{X}_{b,T}^{\mathrm{out}}\in\mathbb{R}^{T\times B\times D}$ as inputs to perform this causal reasoning process.
@@
At each time step, we compute two attention distributions. We first calculate observed attention $A^{\mathrm{obs}}_{\mathrm{in},t}$ using actual exterior features, and then generate counterfactual attention $A^{\mathrm{cf}}_{\mathrm{in},t}$ by replacing all exterior features with their temporal mean $\bar{X}^{\mathrm{out}}=\frac{1}{TB}\sum_{u=1}^{T}\sum_{b=1}^{B}X^{\mathrm{out}}_{u,b}$, which serves as a neutral baseline that removes environmental variations. The difference $\Delta^{\mathrm{in}}_{t}$ between these two distributions quantifies the direct causal influence of external context on interior representations. Specifically, for $t=1,\dots,T$, we obtain
$A^{\mathrm{obs}}_{\mathrm{in},t}$ $=\mathcal{A}\bigl{(}X^{\mathrm{in}}_{t},\,X^{\mathrm{out}}_{\leq t-1},\,X^{\mathrm{out}}_{\leq t-1}\bigr{)},$ (6)
@@
where $\mathcal{A}(Q,K,V)$ denotes multi‐head scaled dot‐product attention. Similarly, $\Delta^{\mathrm{out}}_{t}$ is defined by interchanging the two streams. We further orthogonalize each causal residual against the global baseline vector $\bar{X}$ to ensure that the identified causal patterns reflect true intent-relevant dependencies rather than dataset-specific biases. Formally,
$\Delta_{t}^{\perp}=\Delta_{t}-\frac{\langle\Delta_{t},\,\bar{X}\rangle}{\|\bar{X}\|^{2}+\varepsilon}\,\cdot\bar{X},$ (7)
where $\varepsilon$ is a small constant to ensure numerical stability. The orthogonal projection produces $\Delta^{\perp}_{t}$ that captures context-deviating residuals while removing baseline-aligned components. We retain the final-step residual $\Delta^{\perp}_{T}$ as the decision-relevant signal for downstream modules, ensuring that subsequent fusion operates exclusively on temporally salient causal information.
Dynamic Causal Gating.
@@
Beyond frame-level causal cues, holistic intent understanding requires global semantic reasoning. We extract a coarse intent distribution from the exterior summary $h^{\mathrm{out}}$ through softmax classification over $M$ predefined intention categories ($M$ denotes the total number of classes):
@@
where $W_{\mathrm{int}}\in\mathbb{R}^{M\times D}$ linearly projects the $D$-dimensional exterior summary onto $M$ intention logits. This intention distribution is then re-embedded as a dense intention token $z_{\mathrm{intent}}=W_{\mathrm{proj}}\boldsymbol{\xi}\in\mathbb{R}^{D}$ that encodes the driving intention in continuous representation. The intention token serves as a global semantic anchor, providing top-down guidance across processing streams for consistent interpretation of ambiguous scenarios.
@@
Rong (Rong, Akata, and Kasneci 2020) - - 84.3
@@
The CPE module provides a set of disentangled feature vectors corresponding to internal and external cues and a preliminary intent token. We further introduce the Feature Synthesis Network (FSN), which performs adaptive fusion of these features to construct a superior synthesized representation for predicting driving intention. By selectively emphasizing the most relevant information, the FSN module enhances the robustness of driving intention prediction. Each visual branch undergoes a residual nonlinear transformation via a dual-stage feedforward network with intermediate activation, which, combined with the speed feature $s$, yields the fused representations for the internal, external, and interaction streams:
$r_{\mathrm{in}}$ $=f_{\mathrm{in}}\bigl{(}[h_{\mathrm{in}},\,z_{\mathrm{intent}}]\bigr{)}+h_{\mathrm{in}},$ (10)
$r_{\mathrm{out}}$ $=f_{\mathrm{out}}\bigl{(}[h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{out}},$
$r_{\mathrm{ctx}}$ $=f_{\mathrm{ctx}}\bigl{(}[h_{\mathrm{in}},\,h_{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr{)}+h_{\mathrm{in}}+h_{\mathrm{out}}$
where each $f_{\bullet}$ denotes a dual-stage feedforward mapping comprising two fully connected layers separated by a ReLU activation (FC–ReLU–FC). Let $\mathcal{C}=\{\mathrm{in},\mathrm{out},\mathrm{ctx}\}$. Each refined feature $r_{i}$ ($i\in\mathcal{C}$) is mapped to class logits $\ell_{i}$ and a corresponding confidence weight $w_{i}$, which adaptively controls each branch’s contribution:
@@
$\mathcal{L}=\underbrace{\frac{1}{4}\sum_{i\in\ \mathcal{C}}\mathrm{CE}\bigl{(}\ell_{i},\,y\bigr{)}}_{\text{main loss}}\;+\;\underbrace{\alpha\,\mathrm{CE}\bigl{(}\ell_{\mathrm{intent}},\,y\bigr{)}}_{\text{intention loss}}$ (12)
where $\alpha$ is a tunable weight that governs the contribution of the intention‐prediction term. This unified objective integrates class-imbalance mitigation, multi-view fusion, and intention supervision within a cohesive framework.

### v2

# CaTFormer: Causal Temporal Transformer with Dynamic Contextual Fusion for Driving Intention Prediction

@@ Abstract
Accurate prediction of driving intention is key to enhancing the safety and interactive efficiency of human-machine co-driving systems. It serves as a cornerstone for achieving high-level autonomous driving. However, current approaches remain inadequate for accurately modeling the complex spatiotemporal interdependencies and the unpredictable variability of human driving behavior. To address these challenges, we propose CaTFormer, a causal Temporal Transformer that explicitly models causal interactions between driver behavior and environmental context for robust intention prediction. Specifically, CaTFormer introduces a novel Reciprocal Delayed Fusion (RDF) mechanism for precise temporal alignment of interior and exterior feature streams, a Counterfactual Residual Encoding (CRE) module that systematically eliminates spurious correlations to reveal authentic causal dependencies, and an innovative Feature Synthesis Network (FSN) that adaptively synthesizes these purified representations into coherent temporal representations. Experimental results demonstrate that CaTFormer attains state-of-the-art performance on the Brain4Cars dataset. It effectively captures complex causal temporal dependencies and enhances both the accuracy and transparency of driving intention prediction.
Codehttps://github.com/srwang0506/CaTFormer
@@ Introduction
Driver intention prediction is crucial for autonomous driving systems, as it effectively mitigates risks and enhances driving safety. By forecasting potential outcomes several seconds in advance, the system can proactively alert the driver or initiate evasive maneuvers, significantly improving its safety capabilities.
@@
Initially, intention prediction primarily relied on the extraction and fusion of visual features for basic predictions (Huang et al. 2022). However, advancements in sensor technology have enabled the incorporation of multi-modal information, such as GPS coordinates, vehicle speed, map data, and driver head pose (Hu et al. 2021; Mo et al. 2023; Li et al. 2022; Wu et al. 2023). By leveraging complex models for feature extraction, fusion, and prediction, the performance of driver intention prediction has significantly improved (Sui et al. 2021; Guo et al. 2023; Liu et al. 2023; Gao et al. 2023).
@@
Figure 1: Comparison between previous driving intention prediction methods and ours. (a) is an LSTM-GRU framework processing interior and exterior streams independently before concatenation. (b) is our CaTFormer, a Transformer-based model enabling dynamic causal fusion of dual streams with integrated intention priors. Through joint modeling of global-local dependencies and cross-stream interactions, our approach outperforms existing methods.
@@
Despite advancements in driver intention prediction, the rich multi-modal data remains underutilized. Most existing studies simply concatenate or linearly aggregate this information, as shown in Fig. 1 (a). However, given that the driver controls the vehicle, changes in their state directly influence the vehicle’s driving status, indicating a strong dependency. Therefore, we propose to explicitly model the causal relationship between the driver and the environment, highlighting this causality’s decisive impact on the prediction task, as illustrated in Fig. 1 (b).
@@
Specifically, we adopt a Transformer-based architecture and introduce three sequential components to enhance the causal modeling, encode the driver’s intention, and fuse the multi-dimensional features. First, we introduce a Reciprocal Delayed Fusion (RDF) module that cross-fuses interior and exterior features through a shifting mechanism, explicitly establishing a temporal dependency between the two feature streams. As the fused features may contain considerable causal noise, we further devise a Counterfactual Residual Encoding (CRE) module to filter out such noise to obtain a more explicit causal representation. Finally, we utilize a Feature Synthesis Network (FSN) that employs a gating mechanism to integrate the interior, exterior, and interaction representations, enabling the modeling of both local and global causal structures. Our main contributions are as follows:
@@
  • We propose CaTFormer, an efficient Transformer-based framework for driving intention prediction that embeds causal spatio‑temporal reasoning with adaptive multi‑view fusion in a unified end‑to‑end architecture.
  • Through dual-stream reciprocal delayed fusion, CaTFormer explicitly captures dependencies across interior and exterior streams, isolates genuine causal effects via counterfactual attention subtraction, and adaptively integrates complementary visual cues, effectively enhancing robustness under complex driving conditions.
@@
  • Extensive evaluation on the Brain4Cars dataset demonstrates that CaTFormer demonstrates superior performance in driving intention prediction in both highway and urban scenarios.
@@
In the early stages of driving intention prediction research, studies primarily focused on learning spatiotemporal representations directly from raw video, often employing 3D CNN-LSTM architectures for maneuver prediction (Gebert et al. 2019). While some later work attempted to enrich context by fusing interior and exterior streams via convolutional LSTMs for a more comprehensive decision-making basis (Rong et al. 2020), both approaches struggled with limited capacity for modeling long-range, non-consecutive dependencies.
@@
Inspired by human cognitive processes, TIFN (Guo et al. 2023) introduced a state update unit (STU) to integrate environmental context into driver state modeling and extract semantic segmentation features as attention cues. Similarly, another study framed intention prediction as a sequence-labeling task, combining bidirectional LSTMs with a conditional random field to capture the contextual dependencies of driving behaviors (Zhou et al. 2021). Although these methods incorporate multi-source information, they typically learn inter-modal feature correlations implicitly, lacking explicit disentanglement and reasoning of their interactions. Moreover, to enhance model generalization, existing studies have developed personalized prediction models using techniques like domain-adversarial RNN (Tonutti et al. 2019), inverse reinforcement learning (Liu et al. 2025), and a federated learning framework (Zhu et al. 2024).
@@
The Transformer architecture excels at capturing long-range dependencies and global interaction information in temporal data through its unique attention mechanism (Liu et al. 2024). Building on this, CemFormer (Ma et al. 2023) integrates data from both interior and exterior cameras, learning a unified cross-view representation via a spatio-temporal Transformer to infer driver intention directly from their behavior. However, these methods primarily rely on the Transformer’s inherent structure to interpret dependencies. Meanwhile, a non-autoregressive Transformer with hybrid attention has been employed to simultaneously capture the temporal dynamics of a single vehicle and interactions among multiple vehicles (Jiang et al. 2024). DriveTransformer (Jia et al. 2025) further establishes a unified end-to-end framework to handle perception, prediction, and planning tasks in parallel. While this approach treats intention prediction as an integrated component of a scalable system, it may dilute the model’s focus on specific driver-intention features.
@@
Besides, some early works also integrated causal inference methods into intention prediction. For instance, one study built a causal model to capture temporal relationships within invariant representations from driving data, aiming for domain generalization (Hu et al. 2022). Another adopted a driver-centric approach, framing risk object identification as a causal inference problem and introducing a two-stage causal framework (Li et al. 2020). However, these methods typically cover a relatively limited scope of scenarios and conditions.
@@
Building on prior work, we propose a dual‑stream Transformer architecture that leverages a learned intention embedding to explicitly capture both local and global causal dependencies between in‑cabin and external modalities. Through systematic fusion of these multi‑dimensional feature representations, the model achieves highly accurate intention prediction in complex driving environments.
@@
Figure 2: Overview of the CaTFormer pipeline. After data preprocessing, exterior optical flow is encoded by ResNet-18 and interior images by MobileFaceNet to produce dual-stream feature sequences. These are then fed into three core modules: (1) Reciprocal Delayed Fusion (RDF) for temporal feature integration; (2) Counterfactual Residual Encoding (CRE) for causal enhancement and intention embedding; and (3) Feature Synthesis Network (FSN) for dynamic fusion of complementary interior, exterior and interaction views to yield the final driving intention prediction.
@@
As illustrated in Fig. 2, CaTFormer processes a bi‐stream image sequence $\mathcal{I}=\bigl\{(I^{\mathrm{out}}_{b,t},I^{\mathrm{in}}_{b,t})\bigr\}_{b=1,t=1}^{B,T}$ of $B$ synchronized frame pairs over $T$ time steps, where $b$ and $t$ index the sample and temporal frame, respectively. Feature extractors encode each stream into features $F^{\mathrm{out}},F^{\mathrm{in}}\in\mathbb{R}^{B\times T\times D}$, where $D$ denotes the dimensionality of each encoded feature vector. The features pass through a Reciprocal Delayed Fusion (RDF) module for temporal causality modeling via our proposed bi-stream attention, followed by a Counterfactual Residual Encoding (CRE) module to inject learnable causal representations. Finally, visual features ($z_{\mathrm{in}},z_{\mathrm{out}}$ and interaction features $z_{\mathrm{ctx}}$) are adaptively fused by a Feature Synthesis Network (FSN) into a joint prediction $\ell_{\mathrm{joint}}$.
Reciprocal Delayed Fusion (RDF)
On multi‐lane highways, cameras concurrently record the exterior traffic scene and the driver’s interior state. Motivated by bidirectional dependencies between environmental context and driving behavior, our dual‐stream architecture explicitly models their interactions by jointly processing both feature streams.
To model inter-frame temporal precedence, we introduce a temporal delay mechanism in the Key and Value sequences. Specifically, at time step $t$, the attention mechanism accesses only information from the preceding frame t1t-1. Concretely, we define the delayed feature
$\hat{F}_{b,t}=F_{b,t-1}\,\mathbf{1}_{\{t>1\}},\quad\mathbf{1}_{\{t>1\}}=\begin{cases}1,&t>1,\\ 0,&t=1.\end{cases}$ (1)
@@
Under a strict single-frame delay constraint, BDA enriches each frame’s representation by fusing interior and exterior contexts from the immediately preceding timestep. The current interior and exterior features attend bidirectionally to their one-frame delayed counterparts, capturing both temporal coherence and cross-stream coupling. To model diverse associations efficiently, we project into $H$ parallel attention heads (in our experiments, $H=8$) and aggregate their outputs through concatenation and a final linear mapping:
@@
where $d_{k}$ is the key dimension, $\cdot]_{i=1}^{H}$ denotes concatenation across heads, and $W^{O}$ restores the original feature size. Fig. 3 shows the bidirectional query-key-value fusion, highlighting how interior and exterior streams are jointly updated.
Figure 3: Illustration of Bidirectional Dependency Attention (BDA), where buffered interior and exterior features cross-attend to enhance current-frame representations.
@@
Although BDA generates a fused representation $H_{b,t}$ at each spatial-temporal point, some channels may still carry noise or irrelevant signals. Thus, we apply two channel-wise gating layers to adaptively enhance informative features and suppress spurious ones:
$g_{b,t}$ $=\sigma\bigl(W_{2}\,(\mathrm{ReLU}(W_{1}H_{b,t}+b_{1}))+b_{2}\bigr),$ (3)
$\widetilde{H}_{b,t}$ $=g_{b,t}\,\odot\,H_{b,t},$
where $\sigma(\cdot)$ denotes the sigmoid activation function and $\odot$ denotes the Hadamard product between vectors.
@@
$R(x)=x\;\odot\;\frac{s}{\sqrt{\frac{1}{D}\sum_{d=1}^{D}x_{d}^{2}+\epsilon}},$ (4)
${X}_{b,t}=\mathrm{Dropout}\bigl(R(\widetilde{H}_{b,t})\bigr),$ (5)
where $R(\cdot)$ denotes the Root-mean-square normalization function, $s$ is a learnable scaling parameter, $\epsilon$ is a small constant to ensure numerical stability, and $D$ is the dimensionality of the feature vector $x$, yielding channel-calibrated representations across both interior and exterior feature streams.
Counterfactual Residual Encoding (CRE)
Conventional intention prediction architectures aggregate heterogeneous interior and exterior cues under an implicit correlation assumption and thus often mistake coincidental patterns for proper decision drivers. To overcome this limitation, our CRE module contrasts observed and counterfactual cross‐stream attentions to disentangle direct causal contributions. We then selectively amplify only those residuals that genuinely influence driving intention, resulting in improved robustness and generalization in safety‐critical scenarios. Specifically, CRE takes the bidirectionally fused interior and exterior features ${X}_{b,T}^{\mathrm{in}},{X}_{b,T}^{\mathrm{out}}\in\mathbb{R}^{T\times B\times D}$ as inputs to perform this causal reasoning process.
@@
At each time step, we compute two attention distributions. We first calculate observed dependency attention $A^{\mathrm{obs}}_{\mathrm{in},t}$ using actual exterior features, and then generate counterfactual dependency attention $A^{\mathrm{cf}}_{\mathrm{in},t}$ by replacing all exterior features with their temporal mean $\bar{X}^{\mathrm{out}}=\frac{1}{TB}\sum_{u=1}^{T}\sum_{b=1}^{B}X^{\mathrm{out}}_{u,b}$, which serves as a neutral baseline that removes environmental variations. The difference $\Delta^{\mathrm{in}}_{t}$ between these two distributions quantifies the direct causal attention of exterior context on interior representations. Specifically, for $t=1,\dots,T$, we obtain
$A^{\mathrm{obs}}_{\mathrm{in},t}$ $=\mathrm{BDA}\bigl(X^{\mathrm{in}}_{t},\,X^{\mathrm{out}}_{t-1},\,X^{\mathrm{out}}_{t-1}\bigr),$ (6)
@@
where $\mathcal{A}(Q,K,V)$ denotes multi‐head scaled dot‐product attention. Similarly, $\Delta^{\mathrm{out}}_{t}$ is defined by interchanging the two streams. We further orthogonalize each causal residual against the global baseline vector $\bar{X}$ to ensure that the identified causal patterns reflect true intention-relevant dependencies rather than dataset-specific biases. Formally,
$\Delta_{t}^{\perp}=\Delta_{t}-\frac{\Delta_{t}^{\top}\,\bar{X}}{\|\bar{X}\|^{2}+\varepsilon}\,\cdot\bar{X},$ (7)
where $\varepsilon$ ensures numerical stability. The orthogonal projection yields $\Delta^{\perp}_{t}$ by removing baseline-aligned components. We retain $\Delta^{\perp}_{T}$ as the decision-relevant signal, guiding downstream fusion toward temporally salient causal cues.
Dynamic residual gating.
@@
Beyond frame-level causal cues, holistic intention understanding requires global semantic reasoning. We extract a coarse intention distribution from the exterior summary $h^{\mathrm{out}}$ through softmax classification over $M$ predefined intention categories ($M$ denotes the total number of classes):
@@
where $W_{\mathrm{int}}\in\mathbb{R}^{M\times D}$ linearly projects the $D$-dimensional exterior summary onto $M$ logits. This intention distribution is then re-embedded as an intention token $z_{\mathrm{intent}}\in\mathbb{R}^{D}$ that encodes the driving intention in a continuous representation. The intention token serves as a global semantic anchor, providing top-down guidance across processing streams for consistent interpretation of ambiguous scenarios.
@@
Rong (Rong et al. 2020) - - 84.3
@@
The CRE module provides a set of disentangled feature vectors corresponding to interior and exterior cues and a preliminary intention token. We further introduce the Feature Synthesis Network (FSN), which performs adaptive fusion of these features to construct a superior synthesized representation for predicting driving intention. By selectively emphasizing the most relevant information, the FSN module enhances the robustness of driving intention prediction. Each visual branch undergoes a residual nonlinear transformation via a dual-stage feedforward network with intermediate activation, which, combined with the speed feature $s$, yields the fused representations for the interior, exterior, and interaction streams:
$r_{\mathrm{in}}$ $=f_{\mathrm{in}}\bigl([h^{\mathrm{in}},\,z_{\mathrm{intent}}]\bigr)+h^{\mathrm{in}},$ (10)
$r_{\mathrm{out}}$ $=f_{\mathrm{out}}\bigl([h^{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr)+h^{\mathrm{out}},$
$r_{\mathrm{ctx}}$ $=f_{\mathrm{ctx}}\bigl([h^{\mathrm{in}},\,h^{\mathrm{out}},\,z_{\mathrm{intent}},\,s]\bigr)+h^{\mathrm{in}}+h^{\mathrm{out}}$
where each $f_{\cdot}$ denotes a dual-stage feedforward mapping comprising two fully connected layers separated by a ReLU activation (FC–ReLU–FC). Let $\mathcal{C}=\{\mathrm{in},\mathrm{out},\mathrm{ctx}\}$. Each refined feature $r_{i}$ ($i\in\mathcal{C}$) is mapped to class logits $\ell_{i}$ and a corresponding confidence weight $w_{i}$, which adaptively controls each branch’s contribution:
@@
$\mathcal{L}=\underbrace{\frac{1}{4}\sum_{i\in\mathcal{H}}\mathrm{CE}\bigl(\ell_{i},\,y\bigr)}_{\text{main loss}}\;+\;\underbrace{\alpha\,\mathrm{CE}\bigl(\ell_{\mathrm{intent}},\,y\bigr)}_{\text{intention loss}}$ (12)
where $\mathcal{H}=\{\mathrm{in},\mathrm{out},\mathrm{ctx},\mathrm{joint}\}$ denotes the four stream-level heads, and $\alpha$ controls the weight of the intention supervision term. This unified objective integrates class-imbalance mitigation, multi-view fusion, and intention supervision within a cohesive framework.